# On Edge Coloring of Multigraphs

### v4

@@ 3 Extented Tashkinov Tree (ETT)
The following theorem is due to Tashkinov Tashkinov-2000 ; its proof can also be found in Stiebitz et al. StiebSTF-Book (see Theorem 5.1 on page 116).
@@
So a Tashkinov tree with respect to $e$ and $\varphi$ is a tree-sequence obtained from $(y_{0},e,y_{1})$ by TAA, where $y_{0}$ and $y_{1}$ are two ends of the uncolored edge $e$. We point out that, although there might be several ways to construct a closure of $T$ under $\varphi$, the vertex set of these closures is unique.
Let $C$ be an $(\alpha,\beta)$-chain in $\varphi$ and $T$ be a tree-sequence. A proper subpath $P_{ex}$ of $C$ with ends $u$ and $v$ is called a $(\alpha,\beta)$-exit path of $T$ under $\varphi$ if $V(T)\cap V(P_{ex})=\{v\}$ and $\overline{\varphi}(u)\cap\{\alpha,\beta\}\neq\emptyset$; in this case, $v$ is called the corresponding $(\alpha,\beta)$-exit vertex of $T$ and edge $f$ in $\partial(T)\cap P_{ex}$ is called the corresponding $(\alpha,\beta)$-exit edge of $T$. Note that possibly $\overline{\varphi}(v)\cap\{\alpha,\beta\}=\emptyset$, but $C$ must be an $(\alpha,\beta)$-path if it contains an $(\alpha,\beta)$-exit path of $T$. A proper subpath $D$ of $C$ with ends $u$ and $v$ is called a $(\alpha,\beta)$-ear of $T$ if $|V(D)|\geq 3$ and $V(T)\cap V(D)=\{u,v\}$; in this case, $u$ and $v$ are called roots of $D$ in $T$. See Figure 1 for reference.
@@
Consider a $k$-triple $(G,e,\varphi)$. A Tashkinov Series is a tuple $(T_{n+1},S_{n},F_{n},\Theta_{n})$ constructed by the following algorithm, where $T_{n}+f_{n}$ stands for the tree-sequence augmented from $T_{n}$ by adding a boundary edge $f_{n}$ of $T_{n}$. It is worth pointing out that iteration $n$ includes iteration $1$ when $n=1$, but we still write iteration $1$ out for a better presentation.
@@
    If $T_{1}$ has a $(\gamma_{1},\delta_{1})$-ear $Q$ where $\delta_{1}$ is a defective color of $T_{1}$ under $\varphi$ and $\gamma_{1}\in\overline{\varphi}(T_{1})$, we construct $(T_{2},S_{1},F_{1},\Theta_{1})$ as follows; [^2] We shall see later that Theorem 3.1 (A1) and (A2) guarantee the existence of such a $(\gamma_{1},\delta_{1})$-ear $Q$ if $T_{1}$ is not strongly closed under $\varphi$. In fact, for any color $\gamma_{1}\in\overline{\varphi}(T_{1})$, all the $\delta_{1}$ colored defective edges of $T_{1}$ are contained in some $(\gamma_{1},\delta_{1})$-ears except for one which is contained in a $(\gamma_{1},\delta_{1})$-exit path. See more details in the proof of Theorem 3.2. Pick an arbitrary $\delta_{1}$ colored edge $f_{1}\in\partial(T_{1})$ (boundary edge of $T_{1}$) that contains in the $(\gamma_{1},\delta)$-ear $Q$, set $F_{1}=\{f_{1}\}$, $S_{1}=\{\gamma_{1},\delta_{1}\}$, and $\Theta_{1}=IE$ (Here IE stands for Initial Extension which is the extension type). Let $T_{2}$ be a closure of $T_{1}+f_{1}$ under $\varphi$.
@@
    If there is a subscript $h\leq n-1$ with $\Theta_{h}=IE$ and $S_{h}=\{\delta_{h},\gamma_{h}\}$, such that a $(\gamma_{h},\delta_{h})$-ear $O$ of $T_{h}$ contains a vertex outside $V(T_{n})$, we construct $(T_{n+1},S_{n},F_{n},\Theta_{n})$ as follows; Pick an arbitrary edge $f_{n}$ in $O\cap\partial(T_{n})$ such that $O$ contains a path $L$ connecting $f_{n}$ and $V(T_{h})$ with $V(L)\subseteq V(T_{n})$. Set $F_{n}=\{f_{n}\}$, $S_{n}=\{\gamma_{n},\delta_{n}\}$ with $\gamma_{n}=\gamma_{h}$ and $\delta_{n}=\delta_{h}$, and $\Theta_{n}=RE$. Let $T_{n+1}$ be a closure of $T_{n}+f_{n}$ under $\varphi$. This type of extension is called a Revisiting Extension (RE).
@@
    Else, if $T_{n}$ has a $(\gamma_{n},\delta_{n})$-ear $Q$ where $\delta_{n}$ is a defective color of $T_{n}$ under $\varphi$ and $\gamma_{n}\in\overline{\varphi}(T_{n})$, we construct $(T_{n+1},S_{n},F_{n},\Theta_{n})$ as follows; Pick an arbitrary edge $f_{n}\in\partial(T_{n})\cap Q$, set $F_{n}=\{f_{n}\}$, $S_{n}=\{\gamma_{n},\delta_{n}\}$, and $\Theta_{n}=IE$. Let $T_{n+1}$ be a closure of $T_{n}+f_{n}$ under $\varphi$. This type of extension is called an Initial Extension (IE).
@@
In the above algorithm, for each $1\leq i\leq n$, $f_{i},\delta_{i}$, and $\gamma_{i}$ are called the connecting edge, connecting color, and companion color of $T_{i}$, respectively. We say $T_{i}$ is $RE$ finished if for each $h\leq i-1$ with $\Theta_{h}=IE$, all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ are contained in $G[V(T_{i})]$, that is, $T_{i}$ does not satisfy the requirement for an RE extension. So $T_{1}$ is $RE$ finished. For convention, we also say $T_{0}$ is $RE$ finished. It is worth pointing out that since $RE$ extension always has higher priority than $IE$ extension, for an RE extension, there’s only one choice for connecting color and companion color, although there are multiply choices for connecting edges. For each $T_{i}$ and extension type $S_{i-1}$ with $i\geq 1$, we call the index $h$ with $10\leq h\leq i-$ 1 the initial index of $T_{i}$ and $\Theta_{i-1}$, if $h$ is the largest index with $\Theta_{h}=IE$ when $i\geq 2$, and if $h=0$ when $i=1$. So $T_{h}$ is $RE$ finished and $\Theta_{j}=RE$ for each $h<j<i-1$. Define $D_{n}=\cup_{1\leq i\leq n}S_{i}-\overline{\varphi}(T_{n})$. In the remainder of this paper, we reserve notations $T_{i},S_{i},F_{i},\Theta_{i},D_{i},\gamma_{i},$ and $\delta_{i}$ for Tashkinov series build from Algorithm 3.1.
@@
A tree sequence $T$ is called an Extended Tashkinov Tree (ETT) under $\varphi$ if there exists an Tashkinov series $\mathcal{T}$ such that $T\subseteq\mathcal{T}$. $\mathcal{T}$ is called the corresponding Tashkinov series of $T$. The unique nonegative integer $n$ such that $T_{n}\subsetneq T\subseteq T_{n+1}$ is called the rung number of $T$ and denoted by $n(T)$. The sequence $T_{0}\subset T_{1}\subset\dots\subset T_{n}\subset T+f_{n}$ is called the ladder of $T$.
The idea of extending from the ear is natural. In fact, the author first got this idea in 2017 from AsplundMcD16 by Asplund and McDonald, see Theorem 2 and Figure 2 in their paper. But due to difficulties proving elementaryness and keeping the structure stable, gs used a very technical method to extend a Tashkinov tree instead.
Clearly every Tashkinov series is a closed ETT. Note that if $T$ is an ETT then any segment $T^{\prime}$ of $T$ is also an ETT. Let $T$ be an ETT under $\varphi$ and $C$ is a subset of $[k]$. We say that an edge $f$ of $G$ is incident to $T$ if at least one end of $f$ is contained in $T$; this definition applies to edges of $T$ as well. Since our proof consists of a sophisticated sequence of Kempe changes, the concept of stable coloring introduced below will be employed to preserve some important coloring properties of $T$, such as, among others, the color on each edge and the set of colors missing at each vertex. Usually, $C$ is the set of colors assigned to $E(T)$ but not missing at any vertex of $T$.
@@
Let $T_{n}$ be a Tashkinov series with $(n-1)$-rungs under $\varphi$ and $\pi$ be a $(T_{n},D_{n-1},\varphi)$-stable coloring. It is worth pointing out that although $\pi$ is $(T_{n},D_{n-1},\varphi)$-stable, $T_{n}$ might not be an ETT under $\pi$, because a $(\gamma_{i},\delta_{i})$-ear of $T_{i}$ under $\varphi$ could become a $(\gamma_{i},\delta_{i})$-exit path of $T_{i}$ under $\pi$ for some $i\leq n$. Therefore, we introduce the following concept.
@@
[…]
@@
In view of the lemma above, it is easy to see that if $\pi$ is $(T-v,D_{n},\varphi)$-stable, then $\pi$ is also $(T-v,D_{n},\varphi)$-stable, because under $\varphi$, the last edge of $T$ is either colored by a missing color of $V(T-v)$, or it is colored by a color in $D_{n}$. So we have the following corollary.
@@
Let $T$ be an ETT under a $k$-triple $(G,e,\varphi)$ and $\alpha,\beta$ be two colors in $[k]$. We say a cycle or a path $P$ intersects $T$ (resp. a subgrah $H$ of $G$) if $V(P)\cap V(T)\neq\emptyset$ (resp. ($V(P)\cap V(H)\neq\emptyset$)). If there is at most one $(\alpha,\beta)$-path intersecting $T$ under $\varphi$, then we say $\alpha$ and $\beta$ are $T$-interchangeable or interchangeable for $T$. Note that $P_{v}(\alpha,\beta,\varphi)=\{v\}$ is also an $(\alpha,\beta)$-path intersecting $T$ if $\alpha,\beta\in\overline{\varphi}(v)$ for some $v\in V(T)$. Clearly, if $\alpha$ and $\beta$ are $T$-interchangeable all the $(\alpha,\beta)$-chains intersecting $T$ are cycles except for at most one.
@@
A closed ETT $T$ with $n(T)=n$ under $\varphi$ has interchangeability property if any two colors are $T$-interchangeable provided one of them is in $\overline{\varphi}(T)$ when $T$ is $RE$ finished, and if any two colors are $T$-interchangeable provided one of them is in $\overline{\varphi}(T)-\{\gamma_{n}\}$ when $T$ is not $RE$ finished, where $\gamma_{n}=\gamma_{h}$ with $h$ being the initial index of $T$.
@@
A1
: $T$ is elementary under $\varphi$.
A2
: $T_{n+1}$ has the interchangeability property. Consequently, if $T_{n+1}$ is $RE$ finished, then any $(T_{n+1},D_{n},\varphi)$-stable coloring $\varphi^{\prime}$ is a $T_{n+1}$ mod $\varphi$ coloring, and $T_{n+1}$ remains $RE$ finished under $\varphi^{\prime}$.
A2.5
: Suppose $T_{n+1}$ could be extended further to $T_{n+2}$ with $F_{n+1}=\{f_{n+1}\}$, $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$, and initial index $h$. If $\varphi^{\prime}$ is a $(T_{n+1},D_{n+1},\varphi)$-stable coloring such that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$, then $\varphi^{\prime}$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring.
A3
: Assume $T_{n+1}$ is $RE$ finished, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, and $\gamma\in\overline{\varphi}(T_{n+1})$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})\setminus\{\gamma\}$ and $\beta$ is arbitrary. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then $T_{n+1}$ has the same $(\gamma,\delta)$-exit vertex under $\varphi/P$ as $\varphi$.
A3.5:
Assume $T_{n+1}$ could be extended further with $F_{n+1}=\{f_{n+1}\}$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})\setminus\{\gamma_{n+1}\}$ and $\beta$ is arbitrary. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then the coloring $\varphi^{\prime}=\varphi/P$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring.
For the reminder of this session, let us present a proof of Conjecture 1 using Theorem 3.1.
@@
[…]

### v5

@@ 3 Extented Tashkinov Tree (ETT)
The following theorem is due to Tashkinov [14]; its proof can also be found in Stiebitz et al. [13] (see Theorem 5.1 on page 116).
@@
So a Tashkinov tree with respect to $e$ and $\varphi$ is a tree-sequence obtained from $(y_{0},e,y_{1})$ by TAA, where $y_{0}$ and $y_{1}$ are two ends of the uncolored edge $e$. We point out that, although there might be several ways to construct a closure of $T$ under $\varphi$, the vertex set of these closures is unique. Indeed, for two closures $T^{\prime}_{1}=(y_{0},e_{1},y_{1},\ldots,e_{n},y_{n})$ and $T^{\prime}_{2}$ of $T$, if $V(T^{\prime}_{1})-V(T^{\prime}_{2})\neq\emptyset$, then for the smallest $i$ with $y_{i}\in V(T^{\prime}_{1})$ but not in $V(T^{\prime}_{2})$, we have $\varphi(e_{i})\in\overline{\varphi}(T^{\prime}_{2})$ and $\varphi(e_{i})\in\partial(T^{\prime}_{2})$, a contradiction.
Let $C$ be an $(\alpha,\beta)$-chain under $\varphi$ and $T$ be a tree-sequence. A proper subpath $P_{ex}$ of $C$ with ends $u$ and $v$ is called an $(\alpha,\beta)$-exit path of $T$ under $\varphi$ if $V(T)\cap V(P_{ex})=\{v\}$ and $\overline{\varphi}(u)\cap\{\alpha,\beta\}\neq\emptyset$; in this case, $v$ is called the corresponding $(\alpha,\beta)$-exit vertex of $T$ and edge $f$ in $\partial(T)\cap E(P_{ex})$ is called the corresponding $(\alpha,\beta)$-exit edge of $T$. Note that possibly $\overline{\varphi}(v)\cap\{\alpha,\beta\}=\emptyset$, but $C$ must be an $(\alpha,\beta)$-path if it contains an $(\alpha,\beta)$-exit path of $T$. A proper subpath $D$ of $C$ with ends $u$ and $v$ is called a $(\alpha,\beta)$-ear of $T$ if $|V(D)|\geq 3$ and $V(T)\cap V(D)=\{u,v\}$; in this case, $u$ and $v$ are called roots of $D$ in $T$. See Figure 1 for reference.
@@
Consider a $k$-triple $(G,e,\varphi)$, a Tashkinov Series is a series of tuples $(T_{n+1},S_{n},F_{n},\Theta_{n})$ for $n=0,1,2\ldots$ constructed inductively using the following algorithm, where $T_{n}+f_{n}$ stands for the tree-sequence augmented from $T_{n}$ by adding a boundary edge $f_{n}$ of $T_{n}$. It is worth pointing out that iteration $n$ includes iteration $1$ when $n=1$, but we still write iteration $1$ out for a better presentation.
@@
    If $T_{1}$ has a $(\gamma_{1},\delta_{1})$-ear $Q$ where $\delta_{1}$ is a defective color of $T_{1}$ under $\varphi$ and $\gamma_{1}\in\overline{\varphi}(T_{1})$, we construct $(T_{2},S_{1},F_{1},\Theta_{1})$ as follows; [^2] We shall see later that Theorem 3.2 (A1) and (A2) guarantee the existence of such a $(\gamma_{1},\delta_{1})$-ear $Q$ if $T_{1}$ is not strongly closed under $\varphi$. In fact, for any color $\gamma_{1}\in\overline{\varphi}(T_{1})$, all the $\delta_{1}$ colored defective edges of $T_{1}$ are contained in some $(\gamma_{1},\delta_{1})$-ears except for one which is contained in a $(\gamma_{1},\delta_{1})$-exit path. See more details in the proof of Theorem 3.3. Pick an arbitrary $\delta_{1}$ colored edge $f_{1}\in\partial(T_{1})$ (boundary edge of $T_{1}$) that is contained in the $(\gamma_{1},\delta)$-ear $Q$, set $F_{1}=\{f_{1}\}$, $S_{1}=\{\gamma_{1},\delta_{1}\}$, and $\Theta_{1}=\operatorname{IE}$ (Here $\operatorname{IE}$ stands for Initial Extension which is the extension type). Let $T_{2}$ be a closure of $T_{1}+f_{1}$ under $\varphi$.
@@
    If there exists a subscript $h\leq n-1$ such that $\Theta_{h}=\operatorname{IE}$ and $S_{h}=\{\delta_{h},\gamma_{h}\}$, and there is a $(\gamma_{h},\delta_{h})$-ear $O$ of $T_{h}$ containing a vertex outside $V(T_{n})$, we construct $(T_{n+1},S_{n},F_{n},\Theta_{n})$ as follows; Pick an arbitrary edge $f_{n}$ in $O\cap\partial(T_{n})$ such that $O$ contains a path $L$ connecting $f_{n}$ and $V(T_{h})$ with $V(L)\subseteq V(T_{n})$. Set $F_{n}=\{f_{n}\}$, $S_{n}=\{\gamma_{n},\delta_{n}\}$ with $\gamma_{n}=\gamma_{h}$ and $\delta_{n}=\delta_{h}$, and $\Theta_{n}=\operatorname{RE}$. Let $T_{n+1}$ be a closure of $T_{n}+f_{n}$ under $\varphi$. This type of extension is called a Revisiting Extension ($\operatorname{RE}$).
@@
    Else, if $T_{n}$ has a $(\gamma_{n},\delta_{n})$-ear $Q$ where $\delta_{n}$ is a defective color of $T_{n}$ under $\varphi$ and $\gamma_{n}\in\overline{\varphi}(T_{n})$, we construct $(T_{n+1},S_{n},F_{n},\Theta_{n})$ as follows; Pick an arbitrary edge $f_{n}\in\partial(T_{n})\cap Q$, set $F_{n}=\{f_{n}\}$, $S_{n}=\{\gamma_{n},\delta_{n}\}$, and $\Theta_{n}=\operatorname{IE}$. Let $T_{n+1}$ be a closure of $T_{n}+f_{n}$ under $\varphi$. This type of extension is called an Initial Extension ($\operatorname{IE}$).
@@
In the above algorithm, for each $1\leq i\leq n$, $f_{i},\delta_{i}$, and $\gamma_{i}$ are called the connecting edge, connecting color, and companion color of $T_{i}$, respectively. We say $T_{i}$ is $\operatorname{RE}$ finished if for each $h\leq i-1$ with $\Theta_{h}=\operatorname{IE}$, all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ are contained in $G[V(T_{i})]$, that is, $T_{i}$ does not satisfy the requirement for an $\operatorname{RE}$ extension. So $T_{1}$ is $\operatorname{RE}$ finished. For convention, we also say $T_{0}$ is $\operatorname{RE}$ finished.
@@
It is worth pointing out that since $RE$ extensions always have higher priority than $\operatorname{IE}$ extensions, for each $T_{i}$ which we will perform an $\operatorname{IE}$ extension on, $T_{i}$ must be $\operatorname{RE}$ finished. In fact, each time after an $\operatorname{IE}$ extension, we perform a series of $\operatorname{RE}$ extensions until it is $\operatorname{RE}$ finished, before performing the next $\operatorname{IE}$ extension. This mechanism ensures that, before the next $\operatorname{IE}$ extension, all $(\gamma_{n},\delta_{n})$-ears of $T_{n}$ are eventually contained after an $\operatorname{IE}$ extension of $T_{n}$.
For each $T_{i}$ and extension type $\Theta_{i-1}$ with $i\geq 1$, we call the index $h$ with $0\leq h\leq i-1$ the initial index of $T_{i}$ and $\Theta_{i-1}$, if $h$ is the largest index with $\Theta_{h}=\operatorname{IE}$ when $i\geq 2$, and if $h=0$ when $i=1$. So $T_{h}$ is $\operatorname{RE}$ finished and $\Theta_{j}=\operatorname{RE}$ for each h<j<i1h<j<i-1. Define $D_{n}=\cup_{1\leq i\leq n}S_{i}-\overline{\varphi}(T_{n})$. In the remainder of this paper, we reserve notations $T_{i},S_{i},F_{i},\Theta_{i},D_{i},\gamma_{i},$ and $\delta_{i}$ for Tashkinov series build from Algorithm 3.1.
@@
Let $(G,e,\varphi)$ be a $k$-triple and let $\mathcal{T}=\{(T_{i},S_{i-1},F_{i-1},\Theta_{i-1}):0\leq i\leq n+1\}$ be a Tashkinov series constructed from $(G,e,\varphi)$ following Algorithm 3.1. A tree-sequence $T$ is called an extended Tashkinov tree (ETT) constructed from $\mathcal{T}$ if $_{n}\subset T\subseteq T_{n+1}. 𝒯\mathcal{T}$ is called the corresponding Tashkinov series of $T$. The unique nonegative integer $n$ such that $T_{n}\subsetneq T\subseteq T_{n+1}$ is called the rung number of $T$ and denoted by $n(T)$. The sequence $T_{0}\subset T_{1}\subset\dots\subset T_{n}\subset T+f_{n}$ is called the ladder of $T$.
The idea of extending from the ear is natural. In fact, the author first got this idea in 2017 from [2] by Asplund and McDonald, see Theorem 2 and Figure 2 in their paper. But due to difficulties proving elementaryness and keeping the structure stable, [6] used a very technical method to extend a Tashkinov tree instead.
Clearly every Tashkinov series is a closed ETT. Note that if $T$ is an ETT then any segment $T^{\prime}$ of $T$ is also an ETT. Let $T$ be an ETT under $\varphi$ and $C$ be a subset of $[k]$. We say that an edge $f$ of $G$ is incident to $T$ if at least one end of $f$ is contained in $T$; this definition applies to edges of $T$ as well. Since our proof consists of a sophisticated sequence of Kempe changes, the concept of stable coloring introduced below will be employed to preserve some important coloring properties of $T$, such as, among others, the color on each edge and the set of colors missing at each vertex. Usually, $C$ is the set of colors assigned to $E(T)$ but not missing at any vertex of $T$.
@@
Let $T_{n}$ be a Tashkinov series with (n1)(n-1)-rungs under $\varphi$ and $\pi$ be a $(T_{n},D_{n-1},\varphi)$-stable coloring. It is worth pointing out that although $\pi$ is $(T_{n},D_{n-1},\varphi)$-stable, $T_{n}$ might not be an ETT under $\pi$, because a $(\gamma_{i},\delta_{i})$-ear of $T_{i}$ under $\varphi$ could become a $(\gamma_{i},\delta_{i})$-exit path of $T_{i}$ under $\pi$ for some $i\leq n$. Therefore, we introduce the following concept.
@@
Let $\mathcal{T}=\{(T_{i},S_{i-1},F_{i-1},\Theta_{i-1})\ :\ 1\leq i\leq n+1\}$ be a Tashinov series. A $(T_{n+1},D_{n},\varphi)$-stable coloring $\pi$ is called $T_{n+1}$ mod $\varphi$ if $\mathcal{T}=\{(T_{i},S_{i-1},F_{i-1},\Theta_{i-1})\ :\ 1\leq i\leq n\}$ remains a Tashinov series under $\pi$; and a $(T_{n},D_{n},\varphi)$-stable coloring $\phi$ is called $T_{n}+f_{n}$ mod $\varphi$ if $\mathcal{T}=\{(T^{*}_{i},S_{i-1},F_{i-1},\Theta_{i-1})\ :\ 1\leq i\leq n\}$ is a Tashkinov series under $\phi$, where $T^{*}_{i}=T_{i}$ for each $1\leq i\leq n$, and $T^{*}_{n+1}$ is a closure of $T_{n}+f_{n}$ under $\phi$. For convention, a $T_{0}+f_{0}$ mod $\varphi$ coloring is defined as any $k$-edge-coloring of GeG-e, $T_{0}+f_{0}$ is just a tashkinov tree with the uncolored edge $e$ and its two ends, and a closure of $T_{0}+f_{0}$ is a closure of $e$.
Let $T$ be an ETT under $\varphi$ with $n$-rungs and corresponding Tashkinov series $\mathcal{T}=\{(T_{i},S_{i-1},$ $F_{i-1},\Theta_{i-1})\ :\ 1\leq i\leq n+1\}$. Clearly, if $\pi$ is both $(T,D_{n},\varphi)$-stable and $\varphi$ mod $T_{n}+f_{n}$, then $T$ or any tree-sequence obtained from $T_{n}+f_{n}$ by TAA under $\pi$ is also an ETT with the same ladder, connecting edges, connecting colors, companion colors, and extension types as under $\varphi$. Moreover, mod colorings form equivalent classes. Let $v$ be the last vertex of $T$. Denote by TvT-v the segment of $T$ ending at the vertex before $v$ along $\prec$. Let $T$ be an ETT with $n$-rungs and $T_{n}\subset T$ under $\varphi$. We call a $(T_{n},D_{n},\varphi)$-stable coloring $\pi$ $(T,D_{n},\varphi)$-weakly stable if $\varphi(f)=\pi(f)$ for each $f\in E(T)$ and $\overline{\varphi}(u)=\overline{\pi}(u)$ for each $u\in V(T-v)$.
@@
In view of the lemma above, it is easy to see that if $\pi$ is $(T-v,D_{n},\varphi)$-stable, then $\pi$ is also $(T-v,D_{n},\varphi)$-stable, because under $\varphi$, the last edge of $T$ is either colored by a missing color of V(Tv)V(T-v), or it is colored by a color in $D_{n}$. So we have the following corollary.
@@
Let $T$ be an ETT under a $k$-triple $(G,e,\varphi)$ and $\alpha,\beta$ be two colors in $[k]$. We say a cycle or a path $P$ intersects $T$ (resp. a subgrah $H$ of $G$) if $V(P)\cap V(T)\neq\emptyset$ (resp. ($V(P)\cap V(H)\neq\emptyset$)). If there is at most one $(\alpha,\beta)$-path intersecting $T$ under $\varphi$, then we say $\alpha$ and $\beta$ are $T$-interchangeable or interchangeable for $T$. Note that $P_{v}(\alpha,\beta,\varphi)=\{v\}$ is also an $(\alpha,\beta)$-path intersecting $T$ if $\alpha,\beta\in\overline{\varphi}(v)$ for some $v\in V(T)$. Clearly, if $\alpha$ and $\beta$ are $T$-interchangeable, then all the $(\alpha,\beta)$-chains intersecting $T$ are cycles except for at most one.
@@
A closed ETT $T$ with $n(T)=n$ under $\varphi$ has interchangeability property if any two colors are $T$-interchangeable provided one of them is in $\overline{\varphi}(T)$ when $T$ is $\operatorname{RE}$ finished, and if any two colors are $T$-interchangeable provided one of them is in $\overline{\varphi}(T)-\{\gamma_{n}\}$ when $T$ is not $\operatorname{RE}$ finished, where $\gamma_{n}=\gamma_{h}$ with $h$ being the initial index of $T$.
@@
A1: $T$ is elementary under $\varphi$.
A2: $T_{n+1}$ has the interchangeability property. Consequently, if $T_{n+1}$ is $\operatorname{RE}$ finished, then any $(T_{n+1},D_{n},\varphi)$-stable coloring $\varphi^{\prime}$ is a $T_{n+1}$ mod $\varphi$ coloring, and $T_{n+1}$ remains $\operatorname{RE}$ finished under $\varphi^{\prime}$.
A2.5: Suppose $T_{n+1}$ could be extended further to $T_{n+2}$ with $F_{n+1}=\{f_{n+1}\}$, $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$, and initial index $h$. If $\varphi^{\prime}$ is a $(T_{n+1},D_{n+1},\varphi)$-stable coloring such that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$, then $\varphi^{\prime}$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring.
A3: Assume $T_{n+1}$ is $\operatorname{RE}$ finished, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, and $\gamma\in\overline{\varphi}(T_{n+1})$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})\setminus\{\gamma\}$ and $\beta$ is arbitrary. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then $T_{n+1}$ has the same $(\gamma,\delta)$-exit vertex under $\varphi/P$ as under $\varphi$.
A3.5: Assume $T_{n+1}$ could be extended further with $F_{n+1}=\{f_{n+1}\}$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})\setminus\{\gamma_{n+1}\}$ and $\beta$ is arbitrary. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then the coloring $\varphi^{\prime}=\varphi/P$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring.
For the remainder of this session, let us present a proof of Conjecture 1 using Theorem 3.2.
@@
Assume that $\chi^{\prime}(G)\geq\Delta(G)+2$. Then $G$ must contain a $k$-critical subgraph $H$ with $k=\chi^{\prime}(G)-1\geq\Delta(G)+1\geq\Delta(H)+1$. Let $(H,e,\varphi)$ be a $k$-triple, and construct a Tashkinov series $\mathcal{T}=\{(T_{i},S_{i-1},F_{i-1},\Theta_{i-1})\ :\ 1\leq i\leq n+1\}$ with maximum $n$ following Algorithm 3.1. We claim that $T_{n+1}$ is elementary and strongly closed. Indeed, $T_{n+1}$ is elementary by Theorem 3.2 (A1). Moreover, it is closed as it is a closure of $T_{n}+f_{n}$ under $\varphi$. Since $n$ is maximum, we can not keep extending $T_{n+1}$ by an $\operatorname{RE}$ extension, so $T_{n+1}$ is $\operatorname{RE}$ finished. Suppose on the contrary, that $T_{n+1}$ is not strongly closed. Then there exists a defective color $\delta_{n+1}$ of $T_{n+1}$. Then $\delta_{n+1}\notin\overline{\varphi}(T_{n+1})$ because $T_{n+1}$ is closed. Pick an arbitrary color $\gamma_{n+1}\in\overline{\varphi}(T_{n+1})$. By Theorem 3.2 (A2), $\gamma_{n+1}$ and $\delta_{n+1}$ are $T_{n+1}$-interchangeable, so $P=P_{v}(\gamma_{n+1},\delta_{n+1},\varphi)$ is the only $(\gamma_{n+1},\delta_{n+1})$-path intersecting $T_{n+1}$, where $v$ is the only vertex in $T_{n+1}$ with $\gamma_{n+1}\in\overline{\varphi}(v)$. As the end $v$ of $P$ is in $T_{n+1}$, we only have one $(\gamma_{n+1},\delta_{n+1})$-exit vertex and one $(\gamma_{n+1},\delta_{n+1})$-exit edge of $T_{n+1}$. Therefore, a $\delta_{n+1}$ colored boundary edge $f_{n+1}$ other than the $(\gamma_{n+1},\delta_{n+1})$-exit edge, is contained in some $(\gamma_{n+1},\delta_{n+1})$-ear of $T_{n+1}$. Therefore, we could extend $T_{n+1}$ using $f_{n+1}$ with an $\operatorname{IE}$ extension, a contradiction to $n$ being maximum.
So $T_{n+1}$ is elementary and strongly closed. Now for each color $\alpha$ in $\overline{\varphi}(T_{n+1})$, its color class forms a matching for all vertices in $T_{n+1}$ but the one missing the color $\alpha$. So this gives $|\overline{\varphi}(T_{n+1})|\cdot(|V(T_{n+1})|-1)/2$ many edges in $H[V(T_{n+1})]$ and $|V(T_{n+1})|$ is odd. On the other hand, any color $\beta$ in $k]-\overline{\varphi}(T_{n+1})$ appears exactly once on $\partial(T_{n+1})$, so its color class also forms a matching for all vertices in $T_{n+1}$ but the one incident with the $\beta$ colored boundary edge. Recall that $e$ is uncolored. So together we have $k\cdot(|V(T_{n+1})|-1)/2+1$ many edges in $H[V(T_{n+1})]$, and therefore $\Gamma(H)\geq\frac{k\cdot(|V(T_{n+1})|-1)/2+1}{(|(V(T_{n+1})|-1)/2}>k$. Thus $k+1\leq\Gamma(H)$. Since $\Gamma(G)$ is a lower bound for $\chi^{\prime}(G)$, we have $\Gamma(G)\leq\chi^{\prime}(G)=k+1\leq\Gamma(H)\leq\Gamma(G)$, forcing $\chi^{\prime}(G)=\Gamma(G)$. ∎